OpenStudy (anonymous):

3^(x+4)=7^(x-1) help please

OpenStudy (anonymous):

I answered you yesterday.

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

he did?@ how coudl you em :p

OpenStudy (anonymous):

i still dont get it .

OpenStudy (anonymous):

What have you tried after that?

OpenStudy (anonymous):

idk i need help

OpenStudy (anonymous):

what help? you need the answer?

OpenStudy (anonymous):

yeah

OpenStudy (tommo_lcfc):

I'm interested to hear it because I'm stuck too.

OpenStudy (anonymous):

lololols mwhaha

OpenStudy (anonymous):

http://openstudy.com/users/foolformath#/updates/4f9805f2e4b000ae9ecd98ad Yesterday I did up to this \[ x = \frac{ -(\log7+\log 3^4)}{(\log 3 -\log 7)} \] there is only one step remaining i.e the simplified form which is \[ x= \frac{\log (567)}{\log\left(\frac{7}{3}\right)} \]

OpenStudy (tommo_lcfc):

I finally got x to be 7.483 (to 3 dp).